help asap

#13] Evaluate SSxyds where is the part of the Splane 2x-y+ Z = 4 that corresponds to 1= xt1, osyal
Answer #1

② ssauds = say ritzt z da Ä is projection (R) Di onto Xyplane, Given 22-Y+z=4. a =) Z=4-20+4 Zx==2 zy=1 - 431 03,4?) ittity x